1. X-2/2=y-3/-1=z+4/3
  2. X-3=y+1/3=z-1/-2

Find the equation of plane which containing the two lines


1
Expert's answer
2021-05-31T15:33:48-0400

Equation of any plane through the first line is 


"a(x-2)+b(y-3)+c(z+4)=0,"

where


"2a-b+3c=0"

This plane will pass through the second line if a point on the second line "(3, -1, 1)" lies on it i.e., if 

"a(3-2)+b(-1-3)+c(1+4)=0."

  Then


"\begin{matrix}
 2a-b+3c=0 \\
 a-4b+5c=0
\end{matrix}"

"\begin{matrix}
 -7b+7c=0 \\
 a=4b-5c
\end{matrix}"

"\begin{matrix}
 a=-c \\
 b=c
\end{matrix}"

Hence, the euation of required plane is


"-(x-2)+(y-3)+(z+4)=0"




"(x-2)-(y-3)-(z+4)=0"

Or


"x-y-z-3=0"